The television show Ghost Whistler has been successful for many years. That show recently had a share of 25, meaning that among the TV sets in use, 25% were tuned to Ghost Whistler. Assume that an advertiser wants to verify that 25% share value by conducting its own survey, and a pilot survey begins with 14 households have TV sets in use at the time of a Ghost Whistler broadcast.

Find the probability that none of the households are tuned to Ghost Whistler.
P(none) =

Find the probability that at least one household is tuned to Ghost Whistler.
P(at least one) =

Find the probability that at most one household is tuned to Ghost Whistler.
P(at most one) =

If at most one household is tuned to Ghost Whistler, does it appear that the 25% share value is wrong? (Hint: Is the occurrence of at most one household tuned to Ghost Whistler unusual?)

  • no, it is not wrong
  • or
  • yes, it is wrong
Answer #1

X ~ B ( n = 14 , P = 0.25 )

Part a)

Part b)

P ( X >= 1 ) = 1 - P ( X = 0 ) = 1 - 0.0178 = 0.9822

Part c)


P ( X <= 1 ) = 0.1010

Part d)

P ( X <= 1 ) = 0.1010

The probability is > 5% i.e > 0.05, hence it is not unusual.

no, it is not wrong